• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Feuerbach

    Log entry time 16:13:26 on June 16, 2005

    Entry number 146224

    keyword=Plan for this evening

    The magnet power supplies are still under repair, so E94-107
    still cannot take production data. The technical crew will stop
    working at 8pm tonight, at which point the Hall will be swept and
    locked up in preparation for beam.

    So, by 9pm the hall should be ready to receive beam. The
    activities for tonight are:

    9pm Calibration of the SLI (Arne Freyberger) (1hr)
    -> Arne, please call the Counting House when you are
    finished.

    10pm Calibration of the nA cavities (~3hrs)
    (B.Michaels, J.Musson, L.Kaufman)

    ~2am Effects of shielding on Detector rates (4hrs)
    (B.Wojtsekhowski, N.Liyanage)

    6am Beam delivery to the hall will cease, and a request for the
    Hall to be surveyed in preparation to going to Restricted
    Access for the day shift to continue repairs.

    The SWING and OWL shifts are ACTIVE to ensure shift coverage
    during this time.

    If you have any questions, please contact the Run Coordinator at
    867-1787.